OpenStudy (anonymous):

medal for this Show how to solve seven-twelfths times six-elevenths with and without cancellation.

OpenStudy (mathmate):

\[ \frac{7}{12}\times \frac{6}{11} = ? \]

OpenStudy (anonymous):

This is not a "solve", really; it's a "compute". The only difference is order of operations: 7/12 * 6/11 = (7*6) / (12*11) = 42 / 132 = 7 / 22 OR 7/12 * 6/11 = (7*6) / (12*11) = (7*1) / (2*11) = 7 / 22

OpenStudy (calculusxy):

First do the numerator and then do the denominator. That will give you your answer, but if you want to cancel it out, then write the expression for the numerator next to the numerator and the expression for the denominator next to the denominator. Them cancel out the necessary answers in which they share a common factor if they are diagonal.